# CaesarNeRF: Calibrated Semantic Representation for Few-Shot Generalizable Neural Rendering

### v1

@@ Abstract
$*{}^{*}$qual contribution.    ${}^{\dagger}$orresponding author.
Generalizability and few-shot learning are key challenges in Neural Radiance Fields (NeRF), often due to the lack of a holistic understanding in pixel-level rendering. We introduce CaesarNeRF, an end-to-end approach that leverages scene-level CAlibratEd SemAntic Representation along with pixel-level representations to advance few-shot, generalizable neural rendering, facilitating a holistic understanding without compromising high-quality details. CaesarNeRF explicitly models pose differences of reference views to combine scene-level semantic representations, providing a calibrated holistic understanding. This calibration process aligns various viewpoints with precise location and is further enhanced by sequential refinement to capture varying details. Extensive experiments on public datasets, including LLFF, Shiny, mip-NeRF 360, and MVImgNet, show that CaesarNeRF delivers state-of-the-art performance across varying numbers of reference views, proving effective even with a single reference image. The project page of this work can be found here.
Figure 1: Novel view synthesis for novel scenes using ONE reference view on Shiny [59], LLFF [36], and MVImgNet [68] (top to bottom). Each pair of images corresponds to the results from GNT [52] (left) and CaesarNeRF (right).
@@ 1 Introduction
Generalizing NeRF to novel scenes often involves using pixel-level feature embeddings encoded from input images, as seen in existing methods [67, 51]. These methods adapt NeRF to novel scenes by separating the scene representation from the model through an image encoder. However, relying solely on pixel-level features has its drawbacks: it requires highly precise epipolar geometry and often overlooks occlusion in complex scenes. Moreover, employing pixel-level features ignores the inherent interconnections within objects in the scene, treating the prediction of each pixel independently. This becomes problematic with a limited number of input reference images, as the data scarcity amplifies prediction ambiguity, significantly influenced by the biases of the input camera views.
We present CaesarNeRF, a method that advances the generalizability of NeRF by incorporating calibrated semantic representation. This enables rendering from novel viewpoints using as few as one input reference view, as depicted in Figure 1. Our approach combines semantic scene-level representation with per-pixel features, enhancing consistency across different views of the same scene. The encoder-generated scene-level representations capture both semantic features and biases linked to specific camera poses. When reference views are limited, these biases can introduce uncertainty in the rendered images. To counter this, CaesarNeRF integrates camera pose transformations into the semantic representation, hence the term calibrated. By isolating pose-specific information from the scene-level representation, our model harmonizes features across input views, mitigating view-specific biases and, in turn, reducing ambiguity. In addition, CaesarNeRF introduces a sequential refinement process, which equips the model with varying levels of detail needed to enhance the semantic features. Extensive experiments on datasets such as LLFF [36], Shiny [59], mip-NeRF 360 [4], and the newly released MVImgNet [68] demonstrate that CaesarNeRF outperforms current state-of-the-art methods, proving effective in generalizable settings with as few as one reference view.
@@ 2 Related work
Neural Radiance Field (NeRF) implicitly captures the density and appearance of points within a scene or object [38, 37] and enable rendering from novel camera positions. In recent years, NeRF has witnessed improvements in a wide range of applications, such as photo-realistic novel view synthesis for large-scale scenes [34, 60, 71], dynamic scene decomposition and deformation [76, 31, 40, 41, 42, 44, 75, 21, 27], occupancy or depth estimation [62, 58, 74], scene generation and editing [70, 35, 43, 28, 20, 63, 30, 64, 1], and so on. Despite these advances, most methods still rely on the original NeRF and require retraining or fine-tuning for novel scenes not covered in the training data.
Figure 2: Overview of CaesarNeRF. CaesarNeRF employs a shared encoder to capture two types of features from input views, including scene-level semantic representation $\{\bm{S}_{n}\}$ and pixel-level feature representation $\{\bm{F}_{n}\}$. Following calibration and aggregation of $\{\bm{S}_{n}\}$ from various views, we concatenate it with the pixel-level fused feature, processed by the view transformer. Subsequent use of the ray-transformer, coupled with sequential refinement, enables us to render the final RGB values for each pixel in the target view.
@@
While these methods offer reasonable reconstructions with few inputs, they typically still require training or fine-tuning for specific scenes. Moreover, these methods usually require at least three reference images. With fewer than three, view-specific biases lead to ambiguity, complicating the rendering. Diffusion-based [32, 12, 73, 49] and other generative methods [24, 77] have been explored for single-view synthesis or generative rendering, yet they are mostly limited to single-object rendering and generally fall short for complex scenes, which often result in a style change as shown in the supplementary material.
@@ 3 The proposed method
We first outline the general framework of existing generalizable NeRF in Section 3.1. Then, we present our proposed CaesarNeRF, as illustrated in Figure 2. This model integrates elements of semantic representation, calibration, and sequential refinement, detailed in Section 3.2, 3.3, and 3.4, respectively. The training objective is given in Section 3.5.
@@ 3.2 Scene-level semantic representation
In GNT [52], which uses a view transformer to aggregate pxiel-level features into an $L$-dimensional vector $\widetilde{\bm{F}}$, we extend this by concatenating $\widetilde{\bm{F}}$ with ${\bf\it S}$ to construct a global-local embedding ${\bf\it E}$, as formulated by:
@@
Figure 3: An illustration of conflicting semantic meanings from multiple viewpoints of the same object. When observing the cup from distinct angles, the features retain spatial information but are inconsistent in the scene-level semantic understanding.
@@ 3.5 Training objectives
Central loss. To ensure frame-wise calibrated semantic features $\{\widetilde{\bm{S}}_{n}\}_{n=1}^{N}$ are consistent when projected onto the same target view, we introduce a central loss, defined as:
@@
Point-wise perceptual loss. During the rendering of a bath of pixels in a target view, we inpaint the ground-truth image by replacing the corresponding pixels with the predicted ones. Then, a perceptual loss [23] is computed between the inpainted image and the target image to guide the training process at the whole-image level.
@@ 4.1 Experimental setups
For the experimental setups, we begin by describing the datasets used in our experiments. This is followed by implementation details of our proposed method and the baseline methods we employed for comparison.
@@
Input
Method
PSNR ($\uparrow$)
LPIPS ($\downarrow$)
SSIM ($\uparrow$)
1-view
PixelNeRF [67]
9.32
0.898
0.264
GPNR [50]
15.91
0.527
0.400
NeuRay [33]
16.18
0.584
0.393
IBRNet [56]
16.85
0.542
0.507
GNT [52]
16.57 0.500 0.424
Ours
18.31 0.435 0.521
2-view
PixelNeRF [67]
11.23
0.766
0.282
GPNR [50]
18.79
0.380
0.575
NeuRay [33]
17.71
0.336
0.646
GeoNeRF [22]
18.76
0.473
0.500
MatchNeRF [8]
21.08 0.272
0.689
MVSNeRF [6]
19.15
0.336
0.704
IBRNet [56]
21.25
0.333
0.685
GNT [52]
20.88
0.251 0.691
Ours
21.94 0.224 0.736
3-view
PixelNeRF [67]
11.24
0.671
0.486
GPNR [50]
21.57
0.288
0.695
NeuRay [33]
18.26
0.310
0.672
GeoNeRF [22]
23.40
0.246
0.766
MatchNeRF [8]
22.30
0.234
0.731
MVSNeRF [6]
19.84
0.314
0.729
IBRNet [56]
23.00
0.262
0.752
GNT [52]
23.21 0.178 0.782
Ours
23.45 0.176 0.794
@@ 4.2 Results and analysis
Input
Method
PSNR ($\uparrow$)
LPIPS ($\downarrow$)
SSIM ($\uparrow$)
1-view
IBRNet [56]
14.93 0.625 0.401
GNT [52]
15.99 0.548 0.400
Ours
17.57 0.467 0.472
2-view
MatchNeRF [8]
20.28 0.278 0.636
MVSNeRF [6]
17.25
0.416
0.577
IBRNet [56]
18.40
0.400
0.595
GNT [52]
20.42 0.327 0.617
Ours
21.47 0.293 0.652
3-view
MatchNeRF [8]
20.77
0.249
0.672
MVSNeRF [6]
18.55
0.343
0.645
IBRNet [56]
21.96
0.281
0.710
GNT [52]
22.47 0.247 0.720
Ours
22.74 0.241 0.723
@@
Input
Method
PSNR ($\uparrow$)
LPIPS ($\downarrow$)
SSIM ($\uparrow$)
1-view
IBRNet [56]
14.12 0.682 0.283
GNT [52]
13.48 0.630 0.314
Ours
15.20 0.592 0.350
2-view
MatchNeRF [8]
17.00 0.566 0.392
MVSNeRF [6]
14.23
0.681
0.366
IBRNet [56]
16.24
0.618
0.360
GNT [52]
15.21
0.559 0.370
Ours
17.05 0.538 0.403
3-view
MatchNeRF [8]
17.26 0.551 0.407
MVSNeRF [6]
14.29
0.674
0.406
IBRNet [56]
17.70
0.555
0.420
GNT [52]
15.59
0.538
0.395
Ours
17.55 0.512 0.430
@@
PSNR ($\uparrow$)
LPIPS ($\downarrow$)
SSIM ($\uparrow$)
1-view
IBRNet
19.14
0.458
0.595
GNT
22.22
0.433
0.678
Ours w/o C.
23.61
0.371
0.718
Ours
24.28 0.334 0.747
2-view
IBRNet
24.38
0.266
0.818
GNT
26.94
0.236
0.850
Ours w/o C.
26.34
0.274
0.817
Ours
27.34 0.215 0.856
3-view
IBRNet
25.53
0.203
0.858
GNT
27.41
0.206
0.870
Ours w/o C.
27.10
0.228
0.850
Ours
27.82 0.190 0.875
4-view
IBRNet
25.99
0.190
0.867
GNT
27.51
0.197
0.875
Ours w/o C.
27.30
0.210
0.862
Ours
27.92 0.181 0.881
5-view
IBRNet
26.12
0.188
0.867
GNT
27.51
0.194
0.876
Ours w/o C.
27.34
0.203
0.865
Ours
27.92 0.179 0.882
Method
PSNR ($\uparrow$)
LPIPS ($\downarrow$)
SSIM ($\uparrow$)
LLFF [36]
23.27
0.212
0.798
NeRF [38]
26.50
0.250
0.811
NeX [59]
27.26 0.179 0.904
GNT [52]
27.24 0.087 0.889
Ours
27.64 0.081 0.904
@@
Method Input
PSNR ($\uparrow$)
SSIM ($\uparrow$)
LPIPS ($\downarrow$)
MatchNeRF [8]
2-view
20.59
0.775
0.276
3-view
22.43
0.805
0.244
Caesar-MatchNeRF
2-view
21.55
0.782
0.268
3-view
22.98
0.824
0.242
IBRNet [56]
1-view
16.85
0.507
0.542
2-view
21.25
0.685
0.333
3-view
23.00
0.752
0.262
Caesar-IBRNet
1-view
17.76
0.543
0.500
2-view
22.39
0.740
0.275
3-view
23.67
0.772
0.242
Table 6: Results on LLFF for few-shot generalization after adapting Caesar to other baseline methods.
@@
${\bf\it R}$ len.
Seq.
Cali.
(Baseline GNT)
20.93
0.185
0.731
Ext.
20.85
0.173
0.735
+32
21.43
0.152
0.763
+64
21.49
0.149
0.766
+96
21.46
0.150
0.766
+128
21.49
0.147
0.763
+ 96
21.53
0.146
0.770
+ 96
21.51
0.147
0.769
+ 96
21.67 0.139 0.781
Table 7: Ablations on the semantic representation length $\bm{R}$, sequential refinement (Seq.) and calibration (Cali.). ‘Ext.’ denotes the extension of per-pixel representation to a length of 64 in GNT.
IBRNet [56]
GPNR [50]
NeuRay [33]
CaesarNeRF (Ours)
Ground-truth
@@
Per-scene optimization. Beyond the multi-scene generalizable setting, we demonstrate per-scene optimization results in Table 5. Testing across 8 categories from the LLFF dataset [36], we calculate the average performance over these scenes. CaesarNeRF consistently outperforms nearly all state-of-the-art methods in the comparison, across all three metrics, showing a significant improvement over our baseline method, GNT [52].
Adaptability. To test the adaptability of our Caesar pipeline, we apply it to two other state-of-the-art methods that use view transformers, namely MatchNeRF [8] and IBRNet [56]. We demonstrate in Table 6 that our enhancements in scene-level semantic understanding significantly boost the performance of these methods across all metrics. This indicates that the Caesar framework is not only beneficial in our CaesarNeRF, which is based on GNT [52], but can also be a versatile addition to other NeRF pipelines.
@@
Visualizations. We present our visualization results in Figure 5, where we compare our method with others using one or two views from the LLFF dataset. Additional visual comparisons are provided in the supplementary materials. These visualizations highlight that in scenarios with few views, our method significantly surpasses the competitors, particularly excelling when only a single view is available. In such cases, CaesarNeRF demonstrates enhanced clarity, with sharper boundaries and more distinct objects.

### v2

@@ Abstract
Generalizability and few-shot learning are key challenges in Neural Radiance Fields (NeRF), often due to the lack of a holistic understanding in pixel-level rendering. We introduce CaesarNeRF, an end-to-end approach that leverages scene-level CAlibratEd SemAntic Representation along with pixel-level representations to advance few-shot, generalizable neural rendering, facilitating a holistic understanding without compromising high-quality details. CaesarNeRF explicitly models pose differences of reference views to combine scene-level semantic representations, providing a calibrated holistic understanding. This calibration process aligns various viewpoints with precise location and is further enhanced by sequential refinement to capture varying details. Extensive experiments on public datasets, including LLFF, Shiny, mip-NeRF 360, and MVImgNet, show that CaesarNeRF delivers state-of-the-art performance across varying numbers of reference views, proving effective even with a single reference image. Equal contribution.    Corresponding author.   This work was done when Haidong Zhu was an intern at Microsoft.
Keywords:
Few-view rendering Generalizable NeRF Neural rendering
Figure 1: Novel view synthesis for novel scenes using ONE reference view on Shiny [67], LLFF [42], and MVImgNet [77] (top to bottom). Each triplet of images corresponds to the results from GNT [60] (left), CaesarNeRF (middle) and groundtruth (right).
@@ 1 Introduction
Generalizing NeRF to novel scenes often involves using pixel-level feature embeddings encoded from input images, as seen in existing methods [76, 59]. These methods adapt NeRF to novel scenes by separating the scene representation from the model through an image encoder. However, relying solely on pixel-level features has its drawbacks: it requires highly precise epipolar geometry and often overlooks occlusion in complex scenes. Moreover, employing pixel-level features ignores the inherent interconnections within objects in the scene, treating the prediction of each pixel independently. Prior attempts to utilize scene-level representation either suffers from style mismatches [35] during scene-wide rendering or have been limited to specific object categories [23, 69, 39]. With few input reference under generalizable settings, these limitations become more pronounced, exacerbating the ambiguity in predictions due to biases from camera viewpoints. Although recent diffusion-based models [56, 18] attempt to address these issues through generative approaches, they struggle to effectively use the input views as contextual references for specific scenes.
We present CaesarNeRF, a method that advances the generalizability of NeRF by incorporating calibrated semantic representation. This enables rendering from novel viewpoints using as few as one input reference view, as depicted in Figure 1. Our approach combines semantic scene-level representation with per-pixel features, enhancing consistency across different views of the same scene. The encoder-generated scene-level representations capture both semantic features and biases linked to specific camera poses. When reference views are limited, these biases can introduce uncertainty in the rendered images. To counter this, CaesarNeRF integrates camera pose transformations into the semantic representation, hence the term calibrated. By isolating pose-specific information from the scene-level representation, our model harmonizes features across input views, mitigating view-specific biases and, in turn, reducing ambiguity. In addition, CaesarNeRF introduces a sequential refinement process, which equips the model with varying levels of detail needed to enhance the semantic features. Extensive experiments on datasets such as LLFF [42], Shiny [67], mip-NeRF 360 [4], and the newly released MVImgNet [77] demonstrate that CaesarNeRF outperforms current state-of-the-art methods with limited reference views available, proving effective in generalizable settings with as few as one reference view. The project can be found here.
@@ 2 Related work
Neural Radiance Field (NeRF) implicitly captures the density and appearance of points within a scene or object [44, 43] and enables rendering from novel camera positions. In recent years, NeRF has witnessed improvements in a wide range of applications, such as photo-realistic novel view synthesis for large-scale scenes [40, 68, 80], dynamic scene decomposition and deformation [85, 34, 46, 47, 48, 50, 84, 24, 30], occupancy or depth estimation [71, 66, 83], scene generation and editing [79, 41, 49, 31, 21, 72, 33, 73, 1], and so on. Despite these advances, most methods still rely on the original NeRF and require retraining or fine-tuning for novel scenes not covered in the training data.
@@
While these methods offer reasonable reconstructions with few inputs, they typically still require training or fine-tuning for specific scenes. Moreover, these methods usually require at least three reference images. With fewer than three, view-specific biases lead to ambiguity, complicating the rendering. Diffusion-based [35, 12, 82, 57] and other generative methods [27, 86] have been explored for single-view synthesis or generative rendering, yet they are mostly limited to single-object rendering and generally fall short for complex scenes, which often result in a style change.
@@ 3 The proposed method
We first outline the general framework of existing generalizable NeRF in Section 3.1. Then, we present our proposed CaesarNeRF, as illustrated in Figure 2. This model integrates elements of semantic representation, calibration, and sequential refinement, detailed in Section 3.2, 3.3, and 3.4, respectively. The training objective is given in 3.5.
@@ 3.1 NeRF and generalizable NeRF
Figure 2: Overview of CaesarNeRF. CaesarNeRF employs a shared encoder to capture two types of features from input views, including scene-level semantic representation $\{\bm{S}_{n}\}$ and pixel-level feature representation $\{\bm{F}_{n}\}$. We use the same encoder for both the scene-level semantic representation and the pixel-level embeddings. Following calibration and aggregation of $\{\bm{S}_{n}\}$ from various views, we concatenate it with the pixel-level fused feature, processed by the view transformer. Subsequent use of the ray-transformer, coupled with sequential refinement, enables us to render the final RGB values for each pixel in the target view. The output features serve as the input for the next stage, indicated by matching line colors.
@@ 3.2 Scene-level semantic representation
In GNT [60], which uses a view transformer to aggregate pixel-level features into an $L$-dimensional vector $\widetilde{\bm{F}}$, we extend this by concatenating $\widetilde{\bm{F}}$ with ${\bf\it S}$ to construct a global-local embedding ${\bf\it E}$, as formulated by:
@@
Figure 3: An illustration of conflicting semantic meanings from multiple viewpoints of the same object. When observing the cup from distinct angles, the features extracted after pooling retain spatial information but are inconsistent in the scene-level semantic understanding, leading to conflicts across various reference images after aggregation.
@@ 3.5 Training objectives
Central loss. Since we project the view-level features from different input camera poses to the shared target view, we introduce a central loss to ensure frame-wise calibrated semantic features $\{\widetilde{\bm{S}}_{n}\}_{n=1}^{N}$ are consistent when projected onto the same target view, which is defined as:
@@
Point-wise perceptual loss. During the rendering of a batch of pixels in a target view, we inpaint the ground-truth image by replacing the corresponding pixels with the predicted ones. Then, a perceptual loss [26] is computed between the inpainted image and the target image to guide the training process at the whole-image level.
@@ 4 Experiments
In this section, we first discuss our experimental setups and other details required for our experiments in Section 4.1, followed by our results for different experimental settings, along with ablation studies and analysis in Section 4.2.
@@ 4.1 Experimental setups
Method 1 reference view 2 reference views 3 reference views
PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$)
PixelNeRF [76] 9.32 0.898 0.264 11.23 0.766 0.282 11.24 0.671 0.486
GPNR [58] 15.91 0.527 0.400 18.79 0.380 0.575 21.57 0.288 0.695
NeuRay [38] 16.18 0.584 0.393 17.71 0.336 0.646 18.26 0.310 0.672
GeoNeRF [25] - - - 18.76 0.473 0.500 23.40 0.246 0.766
MatchNeRF [8] - - - 21.08 0.272 0.689 22.30 0.234 0.731
MVSNeRF [6] - - - 19.15 0.336 0.704 19.84 0.314 0.729
IBRNet [64] 16.85 0.542 0.507 21.25 0.333 0.685 23.00 0.262 0.752
GNT [60] 16.57 0.500 0.424 20.88 0.251 0.691 23.21 0.178 0.782
Ours 18.31 0.435 0.521 21.94 0.224 0.736 23.45 0.176 0.794
@@ 4.2 Results and analysis
Method 1 reference view 2 reference views 3 reference views
PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$)
MatchNeRF [8] - - - 20.28 0.278 0.636 20.77 0.249 0.672
MVSNeRF [6] - - - 17.25 0.416 0.577 18.55 0.343 0.645
IBRNet [64] 14.93 0.625 0.401 18.40 0.400 0.595 21.96 0.281 0.710
GNT [60] 15.99 0.548 0.400 20.42 0.327 0.617 22.47 0.247 0.720
Ours 17.57 0.467 0.472 21.47 0.293 0.652 22.74 0.241 0.723
@@
Method 1 reference view 2 reference views 3 reference views
PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$)
MatchNeRF [8] - - - 17.00 0.566 0.392 17.26 0.551 0.407
MVSNeRF [6] - - - 14.23 0.681 0.366 14.29 0.674 0.406
IBRNet [64] 14.12 0.682 0.283 16.24 0.618 0.360 17.70 0.555 0.420
GNT [60] 13.48 0.630 0.314 15.21 0.559 0.370 15.59 0.538 0.395
Ours 15.20 0.592 0.350 17.05 0.538 0.403 17.55 0.512 0.430
Method 1 reference view 2 reference views 3 reference views 4 reference views 5 reference views
PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$)
IBRNet 19.14 0.458 0.595 24.38 0.266 0.818 25.53 0.203 0.858 25.99 0.190 0.867 26.12 0.188 0.867
GNT 22.22 0.433 0.678 26.94 0.236 0.850 27.41 0.206 0.870 27.51 0.197 0.875 27.51 0.194 0.876
Ours w/o C. 23.61 0.371 0.718 26.34 0.274 0.817 27.10 0.228 0.850 27.30 0.210 0.862 27.34 0.203 0.865
Ours 24.28 0.334 0.747 27.34 0.215 0.856 27.82 0.190 0.875 27.92 0.181 0.881 27.92 0.179 0.882
@@
Method LLFF [42] NeRF [44] NeX [67] GNT [60] Ours
PSNR ($\uparrow$) 23.27 26.50 27.26 27.24 27.64
LPIPS ($\downarrow$) 0.212 0.250 0.179 0.087 0.081
SSIM ($\uparrow$) 0.798 0.811 0.904 0.889 0.904
Table 6: Results on LLFF for few-shot generalization after adapting Caesar to other baselines.
Method 1 reference view 2 reference views 3 reference views
PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
MatchNeRF [8] - - - 20.59 0.775 0.276 22.43 0.805 0.244
 + Caesar - - - 21.55 0.782 0.268 22.98 0.824 0.242
IBRNet [64] 16.85 0.507 0.542 21.25 0.685 0.333 23.00 0.752 0.262
 + Caesar 17.76 0.543 0.500 22.39 0.740 0.275 23.67 0.772 0.242
Table 7: Ablations on the semantic representation length $\bm{R}$, sequential refinement (Seq.) and calibration (Cali.). ‘Ext.’ denotes the extension of per-pixel feature to a length of 64 in GNT.
@@
${\bf\it R}$ len. Seq. Cali.
(Baseline GNT) 20.93 0.185 0.731
Ext. 20.85 0.173 0.735
+32 21.43 0.152 0.763
+64 21.49 0.149 0.766
+96 21.46 0.150 0.766
+128 21.49 0.147 0.763
+ 96 21.53 0.146 0.770
+ 96 21.51 0.147 0.769
+ 96 21.67 0.139 0.781
IBRNet [64] GPNR [58] NeuRay [38] CaesarNeRF Ground-truth
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Target Image GNT CaesarNeRF Target Image GNT CaesarNeRF
Figure 6: Depth estimation prediction using one reference view (first row) and two reference views (second row) as input from LLFF comparing CaesarNeRF with GNT.
Per-scene optimization. Beyond the multi-scene generalizable setting, we demonstrate per-scene optimization results in Table 5. We calculate the average performance over 8 categories from the LLFF dataset [42]. CaesarNeRF consistently outperforms nearly all state-of-the-art methods in the comparison, across all three metrics, showing a significant improvement over our baseline method, GNT [60].
Adaptability. To test the adaptability of our Caesar pipeline, we apply it to two other state-of-the-art methods that use view transformers, namely MatchNeRF [8] and IBRNet [64]. We demonstrate in Table 6 that our enhancements in scene-level semantic understanding significantly boost the performance of these methods across all metrics. This indicates that the Caesar framework is not only beneficial in our CaesarNeRF, which is based on GNT [60], but can also be a versatile addition to other NeRF pipelines with view transformers to aggregate different input views.
@@
Visualizations. We present our visualization results in Figure 5, where we compare our method with others using one or two views from the LLFF dataset. Additional visual comparisons are provided in the supplementary materials. These visualizations highlight that in scenarios with few views, our method significantly surpasses other generalizable NeRF models, particularly excelling when only a single view is available. CaesarNeRF demonstrates rendering with sharper boundaries and more distinct objects.
Depth estimation. We extend our evaluation to depth prediction within the LLFF [42] dataset, focusing on challenges presented by few reference views, such as scenarios with just one or two images. In a comparison between CaesarNeRF and GNT [60], we observe in Figure 6 that GNT struggles to accurately capture the relative positions of objects when reference images are sparse. For instance, with only a single view of a flower, CaesarNeRF precisely indicates the flower’s proximity to the camera compared to the leaves in the background, a distinction that GNT fails to make. Furthermore, the depth estimations provided by CaesarNeRF are consistently more reliable. In the horn example involving two views, CaesarNeRF offers better boundary delineation, showing particular strength in handling reflective surfaces such as grass in the background.
Source Image Smallest Distance $2^{nd}$ Smallest Distance Largest Distance $2^{nd}$ Largest Distance
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
fortress chesstable (0.45) colorfountain (0.50) fenceflower (1.91) pond (1.80)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
fern playground (0.46) bikes (0.50) peppers (2.02) apples (2.01)
Figure 7: Largest and smallest distances for two examples from LLFF test split when matching with training scenes. Numbers (×$\times 10^{-2}$) denote the L-2 distance to the source image within the semantic feature space.
Semantic analysis for $\tilde{{\bf\it S}}$. To assess whether the calibrated semantic representation $\tilde{{\bf\it S}}$ truly captures semantic details of the scene, we analyze the highest and lowest response based on the L-2 distance between features from two different scenes, interpreting smaller distances as greater similarity. We consider the first image from each category in the LLFF dataset as a reference and extract the scene-level representation for the ten closest views to this reference image using CaesarNeRF.
In Figure 7, we present two examples, “room” and “fortress”, from the LLFF dataset, representing the first image of these categories alongside images corresponding to the top-2 highest and lowest responses. This analysis reveals that the scene-level representation predominantly emphasizes structural information and objects of similar categories. For instance, with the source image “room”, the highest responses correlate with images featuring table-like structures, indicating an object-centric focus. In contrast, the lowest responses include images of flowers or ponds found in open areas, which diverge significantly from the structural and object content in the source images.
Comparison with generative methods. Single-view scenarios are often addressed by generative methods [35, 37, 36] that employ diffusion models [19, 53]. While these models can produce reasonable results for object-centric renderings, they struggle with scene-level renderings from novel viewpoints. We show two LLFF examples in Figure 8 using zero 1-to-3 [35], where the left image is the input, and the right one is the output. Images rendered with zero 1-to-3 [35] suffer from style difference. Unlike NeRF-based approaches that reconstruct images from observed pixels, generative models synthesize an entire image from its semantic representation but may not maintain the style.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) leaves (b) orchid
Figure 8: Synthetic results for two examples from LLFF [42], “leaves” and “orchid”, using zero 1-to-3 [35] with one reference image as input and 2 degrees of vertical shift. The left image of each pair is the input, and the right one is the output of zero 1-to-3 [35].